slipknot3066:

What is the fifth term of the sequence? an=5⋅2n−1 Enter your answer in the box. a5=

Shadow:

Do you have an idea of how to start this type of problem off?

slipknot3066:

um no not a clue

Shadow:

Also is it \[a _{n} = 5 \times 2(n - 1)\] or \[a _{n} = 5 \times 2^{(n -1)}\]

slipknot3066:

thanks you saved my life on this

Shadow:

It's the second one?

slipknot3066:

no that was just one question

Shadow:

No I am asking which form is it in?

slipknot3066:

oh the bottom one

Shadow:

Haha, that is an important distinction. If (n-1) is not as an exponent then it starts to look like the form for an arithmetic sequence, except you would have an addition sign instead of multiplication after a1 (or in this case 5)

Shadow:

Arithmetic Sequences: \[a _{n} = a _{1} + d(n - 1)\] Geometric Sequences: \[a _{n} = a _{1} \times r ^{(n - 1)}\] Where d is the difference between two terms and r is the ratio by which the sequence increases. Here, we are dealing with geometric sequences. Since n represents the place number of the term in the sequence, if we are looking for the 5th term of the sequence, we simply input 5 into your given equation of: \[a _{n} = 5 \times 2^{(n -1)}\]

Shadow:

Do you know how to solve for that?

slipknot3066:

um....no

Shadow:

\[a _{5} = 5 \times 2^{(5 - 1)}\] \[a ^{5} = 5 \times 2^{4}\]

Shadow:

Take 2 to the 4th power, then multiply your result by 5

slipknot3066:

so 16 times 5

slipknot3066:

lol 80

Shadow:

Yep :)

slipknot3066:

thanks so much

Shadow:

No problem